OpenStudy (anonymous):

factor completely -2x^2+3x+9 HELP

hartnn (hartnn):

ok, can you find 2 numbers whose sum is +3 and product is -2*9=-18 ?

OpenStudy (anonymous):

i cant im confused

hartnn (hartnn):

what are the factors of 18 ?

OpenStudy (anonymous):

12369

hartnn (hartnn):

yes, so which 2 numbers among then will give you difference of 3 ?

OpenStudy (anonymous):

1and2

hartnn (hartnn):

that will give sum=18 since you have -18, you'll need difference =3

hartnn (hartnn):

*that will give sum=3

OpenStudy (anonymous):

oh ok so 6-3

hartnn (hartnn):

yes! so, you write middle term as 3x =6x-3x

hartnn (hartnn):

-2x^2+3x+9 = -2x^2+6x-3x+9 can you solve further ?

hartnn (hartnn):

just factor -2x from first 2 terms and -3 from other 2.

OpenStudy (anonymous):

so i would get x^2+6x-3 right

hartnn (hartnn):

-2x^2+3x+9 = -2x^2+6x-3x+9 =-2x(....- ....) -3 (...-...) can you fill out the blanks ?

OpenStudy (anonymous):

x-3 x-3

hartnn (hartnn):

yes,-2x^2+3x+9 = -2x^2+6x-3x+9 =-2x(x-3) -3 (x-3) now last step, factor (x-3) from above, what u get ?

OpenStudy (anonymous):

im a lil confused not sure

hartnn (hartnn):

-2x^2+3x+9 = -2x^2+6x-3x+9 =-2x(x-3) -3 (x-3) =(-2x-3)(x-3) its just like ab+cb = (a+c)b

OpenStudy (anonymous):

ok thank you i understand how you get it when i see it complete i just not sure of the operations :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

hope you can help me with future probs ill probably be here all night

hartnn (hartnn):

i hope so, too :) but if i am unavailable, there r many to help you :)